MODERATE RESOLUTION JET COOLED CAVITY RINGDOWN SPECTRA OF THE Ã STATE OF NO_{3} RADICAL

\underline{TERRANCE J. CODD}, MING-WEI CHEN^{a}, MOURAD ROUDJANE and TERRY A. MILLER, Laser Spectroscopy Facility, The Ohio State University, Columbus, Ohio 43210.

The Ã-X̃ spectrum of NO_{3} has been previously observed using cavity ringdown spectroscopy (CRDS) by Andrei Deev et. al under ambient conditions.^{a} There the authors assigned a number of vibronic bands in the spectrum. However, under these conditions, hot-bands may be present and the spectrum becomes very congested at frequencies higher than ∼8700 cm^{−1} due to the density of vibronic states and the overlap of their rotational contours. In order to obtain more information about the Ã state of NO_{3} we recently obtained spectra from 7550 cm^{−1} to over 10000 cm^{−1} using our moderate resolution (≃ 0.05 cm^{−1}) jet cooled CRDS apparatus. Jet cooling in our apparatus reduces the rotational temperature to <30 K and eliminates vibrational hot bands greatly simplifying the spectrum. We are able to resolve and assign more than 15 vibronic features including a new assignment of the 3_{0}^{1} band. Analysis of the ν_{4} progression shows weak Jahn-Teller coupling in this mode. Anomalous band contours and anharmonic spacings are observed for the ν_{1}ν_{4} combination bands and the cause is being investigated. We also see some features that could belong to vibronically forbidden transitions which may be magnetic dipole allowed.

apresent address: University of Illinois at Urbana-Champaign, Urbana, IL 61801

aA. Deev, J. Sommar, and M. Okumura, J. Chem. Phys. 122, 224305 (2005)